# Endoshare: A Source Available Solution to De-Identify and Manage Surgical Videos

### v1

@@ Abstract
Background: Video-based assessment and surgical data science can advance surgical training, research, and quality improvement. However, widespread use is limited by heterogeneous recording formats and privacy concerns associated with video sharing. The present work develops and validates Endoshare, a source available and user-friendly application that merges, standardizes, and de-identifies endoscopic videos guiding minimally invasive surgical procedures.
@@
Results: In the analysis phase, four clinicians and four computer scientists tested a prototype, reporting high usability (4.68 $\pm$ 0.40/5 and 4.03 $\pm$ 0.51/5). The lowest score (4.00 $\pm$ 0.93/5) concerned the clarity of labels, indicating a need to make functions easier to recognize. The application’s user experience was refined to ensure minimal effort for case selection, video merging, automated out-of-body detection and removal, and filename pseudonymization. In the testing phase, ten surgeons reported high perceived usefulness (5.07 $\pm$ 1.75/7), ease of use (5.15 $\pm$ 1.71/7), heuristic usability (4.38 $\pm$ 0.48/5), and recommendation likelihood (9.20 $\pm$ 0.79/10). Benchmarking showed that processing time depended on processing mode, video duration (both $p\leq 0.001$), and machine computational power ($p=0.041$).
Conclusion: Endoshare provides a user-friendly pipeline for standardized, privacy-preserving surgical video management. Endoshare has the potential to be a deployable alternative to proprietary video-management systems, offering transparency, flexibility, and allowing users to retain custody of their data. Compliance certification and broader interoperability validation are needed to establish it as a reliable tool for surgical video management.
@@ 1 Introduction
The objective of this study is to develop and evaluate Endoshare, a novel surgical video management and de-identification application. By releasing this novel solution under a source-available license (PolyForm Noncommercial), the study also aims to contribute significantly toward overcoming current barriers and enabling broader adoption of surgical video collection, sharing, and VBA in clinical and research settings.
2 Methods
This study involved the development and testing of Endoshare. The process followed the first four phases of the software development life cycle (SDLC) [17]: analysis, design, implementation, and testing. Development was iterative and user-centered, incorporating feedback at multiple stages. Two structured surveys were used to guide this process: the first, conducted within the lab but with participants not directly involved in the project, served as a usability evaluation tool during the early stages of development; the second involved external surgeons. A technical assessment was conducted to test the performance of Endoshare on different computing platforms.
2.1 Analysis
An initial internal survey, aligned with established SDLC and usability engineering methodologies [17, 5], was developed during the early development phase to evaluate key system components; this played a central role in the analysis phase to guide further optimization. The questionnaire was constructed using Typeform, and the results were analyzed with Microsoft Excel. Following Nielsen [23], it assessed ten heuristic metrics on a 1–5 scale (Visibility, Match, Control, Consistency, Error, Recognition, Flexibility, Aesthetic, Recover, Help; item descriptions in Table 1), and suggestions were collected for all of them. Participants included both clinicians, to assess usability from an end-user perspective, and computer scientists, to identify workflow and technical optimization needs. Each participant tested the software prototype and provided feedback through both multiple-choice and open-ended questions.
@@ 2.1 Analysis
2.2 Design and Implementation
The application, Endoshare, was designed to process surgical videos. Its development was grounded in principles of transparency and long-term adaptability. This strategy ensures that the system can evolve alongside advancements in video processing, AI-based detection of sensitive visual information , and regulatory requirements. Design decisions emphasized flexibility in both deployment and use. The software architecture was developed to accommodate variable hardware capabilities, institutional constraints, and user expertise levels, from researchers with technical backgrounds to clinicians with minimal computing experience. Dependencies were carefully selected to maximize platform compatibility and minimize installation friction, enabling straightforward setup in typical clinical or academic environments. The project also adopted community-oriented development practices, such as Git-based version control, markdown issue tracking, and standardized code formatting, to maintain clarity and facilitate collaboration. The software is also released under a source-available non-commercial license to ensure that the codebase remains transparent and reproducible, and can be openly reviewed and extended within the academic community.
2.3 Testing
A second survey was conducted to evaluate the fully integrated version of Endoshare in external clinical environments, constituting the user-oriented testing phase of the SDLC. Distributed via Google Forms, the survey targeted practicing surgeons who were given early access to Endoshare, which was installed on their personal or work computers. Building on the heuristic metrics used in the analysis phase, the design incorporated elements from the Technology Acceptance Model (TAM) [27] to assess Perceived Usefulness (PU) and Perceived Ease of Use (PEOU) (Table 1). The survey included both closed and open-ended questions, grouped into the following domains: user demographics and clinical background, prior experience with video processing, system usability (heuristics), evaluation of specific features (merging and de-identification), adherence to usability principles (TAM), assessment of needs for surgical video management, and overall user satisfaction.
@@ 2.3 Testing
2.4 Statistical Analysis
Survey data were analyzed using a mixed-methods approach. Quantitative responses were reported as mean $\pm$ standard deviation (SD). Statistical significance was defined as $p<0.05$ for all analyses. Normality and variance homogeneity were checked via Shapiro–Wilk and Levene’s tests; two-group comparisons used independent-samples t-tests (Student’s or Welch’s, as appropriate). Associations were explored with Pearson/Spearman correlations and multivariable linear regression. Categorical items were examined by Pearson’s $\chi^{2}$ goodness-of-fit. Construct validity was assessed by principal-axis exploratory factor analysis with varimax rotation. Analyses were conducted in Python (SciPy, Statsmodels, Pingouin, FactorAnalyzer, Matplotlib). Qualitative feedback was subjected to inductive thematic analysis by two independent coders. For the technical assessment, mean $\pm$ SD were computed for each Machine $\times$ Mode $\times$ Video condition. Because all machines were tested under every condition, data were analyzed using a linear mixed-effects model with Mode and Video as fixed effects and Machine as a random intercept. Within-video comparisons between fast and advanced modes were assessed using paired t-tests on per-machine means, and geometric mean ratios with bootstrap 95 % confidence intervals were reported as percentage differences. To verify the robustness of these results under potential deviations from normality, a Gamma generalized linear model (GLM) with a log link was fitted including Machine, Mode, and Video as fixed factors. All analyses were conducted in Python using pandas, SciPy, and Statsmodels.
@@ 3 Results
Endoshare is a source available application, freely downloadable from the following website. The source code is available at this GitHub repository. Thorough documentation is provided to support peer review, reproducibility, and future improvements.
3.1 Analysis
The first survey was conducted by four surgeons and four computer scientists. Both groups consistently rated the heuristics highly ($4.68/5\pm 0.40$ vs. $4.03/5\pm 0.51$ for surgeons and computer scientists, respectively; $p=0.096$). Error Prevention was the item that received the highest overall score ($4.63/5\pm 0.52$), while Recognition Rather Than Recall was the lowest ($4.00/5\pm 0.93$).
Qualitative analysis of heuristic-based feedback matched the numeric evaluation of these metrics: it revealed positive remarks regarding system consistency, error prevention, and visual minimalism, with users describing the interface as “easy to follow” and “not distracting from work.” However, comments linked to the heuristics of visibility and recognition highlighted recurring issues with font size, unclear labels, and the lack of visual guidance, suggesting that key elements of the workflow could be made more discoverable and self-explanatory. These items were improved during the development of Endoshare.
3.2 Application
@@ 3.3 Testing
All participants recorded surgical videos primarily for research purposes, with secondary uses including conferences (80%), quality assurance (70%), training (60%), documentation (40%), and education (30%). The most common capture method was via USB drives connected to the endoscopic tower (80%), while alternative methods such as built-in systems, internal server solutions, or cloud upload were used by fewer respondents ($\leq$30%), as reported in Figure 3A. Long-term storage was most frequently on external hard drives (80%), followed by personal computers (30%) and hospital servers or cloud platforms (20%) (Figure 3B).
Analysis revealed statistically significant differences in the distribution of responses for recording method ($\chi^{2}=10.08$, $p=0.018$) and recording setting ($\chi^{2}=7.36$, $p=0.007$), indicating dominant use of tower-based USB recording in the OR. No other variables showed significant differences.
Users reported intention to use Endoshare mainly for clinical and translational research (100%), followed by multicenter study preparation (60%), hospital databases (40%), conferences (30%), and education (10%). Those who will mostly use the platform include researchers (90%), residents or trainees (70%), and surgeons (70%), with less involvement from educators (20%) or nursing/OR staff (20%) (Figure 3C–D).
@@
PU and PEOU were highly correlated ($r=0.969$, $p<0.001$), while heuristic usability correlated with recommendation likelihood ($r=0.727$, $p=0.017$). In contrast, PU ($r=0.284$) and PEOU ($r=0.126$) showed weaker, non-significant associations with recommendation likelihood.
In multiple regression ($R^{2}=0.62$, $p=0.101$), heuristics showed the strongest positive coefficient (+0.88+0.88, $p=0.146$), although no predictor reached statistical significance, possibly due to multicollinearity and small sample size.
Exploratory factor analysis (EFA) extracted three components: PU and PEOU items loaded together on a single factor, while heuristic usability split into two subdomains—one related to system feedback and support, which included recognition, error recovery, help, visibility, consistency, and flexibility; the other related to error prevention and control. The full factor loading matrix is available in the Supplementary Materials (eTable 2).
No statistically significant differences in usability scores were found across subgroups defined by academic qualification, scientific experience, technical familiarity, or frequency of video recording (all $p>0.05$).
With regard to the performance benchmarking, a total of 18 conditions (3 machines $\times$ 2 modes $\times$ 3 video lengths) were tested with $n=3$ repetitions per condition. Mean speed-up times, expressed as a ratio of the video length, are reported in Figure 5.
Fig. 5: Benchmarking Endoshare performance. Processing speed of Endoshare across three hardware configurations (MacOS laptop, Windows average desktop, Linux high-end desktop) for 1-, 30-, and 60-minute videos in both Fast and Advanced modes. Bars indicate mean real-time speed-up (× video length).
A linear mixed-effects model revealed significant main effects of processing mode ($p=0.001$) and video duration ($p<0.001$), but no interaction between them ($p>0.149$). Fast Mode consistently outperformed Advanced Mode, reducing processing time by approximately 74% ($p=0.001$). The effect of video duration was proportional to the expected increase in total processing time, while the relative advantage of Fast Mode remained stable across all video lengths.
Paired comparisons across machines showed that Advanced Mode required markedly longer processing ($3.8\times$, $7.6\times$, and $5.6\times$ higher times for 1-, 30-, and 60-minute videos, respectively), corresponding to an overall $6.2\times$ (+519%) increase in processing time. Although this difference did not reach statistical significance ($p=0.053$), the trend consistently favored the Fast Mode across all durations. Results were consistent under a Gamma generalized linear model ($p=0.001$), confirming robustness to non-normality. Within this model, a likelihood-ratio test for the Machine factor ($\chi^{2}(2)=6.374$, $p=0.041$) indicated a modest but statistically significant contribution of hardware type to overall processing performance. The complete numerical results are provided in the Supplementary Materials (eTable 3).
@@ 4 Discussion
Endoshare, a source available, platform-independent solution, systematically addresses main limitations in surgical video sharing and management. Developed through an iterative, user-centered design approach, Endoshare is an easy to use, surgeon-friendly system to streamline surgical video use, as illustrated in Figure 6.
@@
Fig. 6: Figure 6 Schematic overview of the Endoshare implementation within surgical practice. Endoscopic procedures are recorded and transferred via external storage to any computer with Endoshare installed; the platform then automates merging and de-identification, facilitating use and enabling privacy-preserving sharing of videos for education, multicenter research, and laboratory analysis.
@@
The pipeline starts with the recording of videos guiding minimally invasive procedures; users can then transfer via encrypted external hard drives the videos to any computer equipped with Endoshare. With a single setup step, all videos are automatically prepared for processing and export, making them immediately available for educational or research use. By default, the application launches in “Fast” mode to maximize throughput, while an advanced interface offers power users fine‐grained control over encoding parameters such as resolution and frame rate when needed. A step‐by‐step user guide is bundled with the software to facilitate initial installation and setup.
@@
This workflow was shaped and improved through two distinct evaluation phases. In the analysis phase, the initial usability assessment guided early refinements to Endoshare, particularly in visibility, recognition, and labeling. In the subsequent testing phase, external users confirmed the strengths of the application, especially in error prevention, interface consistency, and minimalistic design, which were reflected in the high heuristic usability scores. In addition, high PU and PEOU scores, together with strong recommendation likelihood, confirmed high user acceptance. Complementing these usability findings, the technical evaluation confirmed Endoshare’s consistent performance across different hardware configurations and video durations. Processing time scaled proportionally with video length, and the system maintained stable operation on all tested machines, demonstrating robustness and cross‐platform compatibility.
@@
Endoshare stands out from commercially available video‐management solutions through its source‐available architecture, minimal hardware dependencies, and ease of deployment. Whereas many proprietary platforms require dedicated servers, vendor‐locked hardware, or extensive IT integration, Endoshare can be easily installed on any Windows, macOS, or Linux workstation and leverages standard file‐system protocols. Its modular design permits rapid inclusion of new codecs or de‐identification algorithms without altering core infrastructure, and its reliance on widely used libraries (PyQt5, FFmpeg, TensorFlow) ensures transparency and reproducibility. In contrast, established clinical video systems often impose licensing fees, rigid workflows, and limited customization, which can hinder adoption—particularly in resource‐constrained or research‐focused settings. By providing a fully auditable codebase and comprehensive user documentation, Endoshare lowers the barrier for centers to implement video sharing while keeping data management fully within the institution.
@@ 4.1 Future Developments
Functionality will also be extended to include manual review and correction of frames misclassified by the automated de‐identification algorithm. A dedicated interface will allow users to inspect flagged scenes, adjust classifications, and reinitiate the de‐identification process as needed. This manual override mechanism is critical for addressing edge cases and improving the overall reliability and trustworthiness of the platform in clinical settings.

### v2

# Endoshare: A Publicly Available, Surgeons-Friendly Solution to De-Identify and Manage Surgical Videos

@@ Abstract
Background: Video-based assessment and surgical data science can advance surgical training, research, and quality improvement. However, widespread use is limited by heterogeneous recording formats and privacy concerns associated with video sharing. The present work develops, evaluates, and publicly releases Endoshare, a surgeon-friendly application that merges, standardizes, and de-identifies endoscopic videos.
@@
Results: In the analysis phase, four clinicians and four computer scientists tested a prototype, reporting high usability (4.68 $\pm$ 0.40/5 and 4.03 $\pm$ 0.51/5). The lowest score (4.00 $\pm$ 0.93/5) concerned the clarity of labels, indicating a need to make functions easier to recognize. The application’s user experience was refined to ensure minimal effort for case selection, video merging, automated out-of-body detection and removal, and filename pseudonymization. In the testing phase, ten surgeons reported high perceived usefulness (5.07 $\pm$ 1.75/7), ease of use (5.15 $\pm$ 1.71/7), heuristic usability (4.38 $\pm$ 0.48/5), and recommendation likelihood (9.20 $\pm$ 0.79/10). Benchmarking showed that processing time increased proportionally with video duration and was consistently lower in fast mode ($\approx$79% faster).
Conclusions: Endoshare is a publicly available surgeon-friendly solution to manage, standardize, and de-identify surgical videos. Endoshare could facilitate the use and sharing of precious data for surgical training, research, and quality improvement. Compliance certification and broader interoperability validation are needed to establish it as a reliable tool for surgical video management.
@@ 1 Introduction
The objective of this study is to develop and evaluate Endoshare, a novel surgical video management and de-identification application. By releasing this novel solution under a source-available license, the study also aims to contribute significantly toward overcoming current barriers and enabling broader adoption of surgical video collection, sharing, and VBA in clinical and research settings.
2 Materials and Methods
This study involved the development and testing of Endoshare. The process followed the first four phases of the software development life cycle (SDLC) [26]: analysis, design, implementation, and testing. Development was iterative and user-centered, incorporating feedback at multiple stages. Two structured surveys were used to guide this process: the first, conducted within the laboratory but with participants not directly involved in the project, served as a usability evaluation tool during the early stages of development; the second involved external surgeons. A technical assessment was conducted to test the performance of Endoshare on different computing platforms.
2.1 Analysis phase
An initial internal survey, aligned with established SDLC and usability engineering methodologies [26, 27], was developed during the early development phase to evaluate key system components; this played a central role in the analysis phase to guide further optimization. The questionnaire was implemented on Typeform, and the results were analyzed with Microsoft Excel. Following Nielsen J. [28], it assessed ten heuristic metrics on a 1–5 scale (Visibility, Match, Control, Consistency, Error, Recognition, Flexibility, Aesthetic, Recover, Help; item descriptions in Table 1), and suggestions were collected for all of them. Participants included both clinicians, to assess usability from an end-user perspective, and computer scientists, to identify workflow and technical optimization needs. Each participant tested the software prototype and provided feedback through both multiple-choice and open-ended questions.
@@ 2.1 Analysis phase
2.2 Design and Implementation phases
The application, Endoshare, was designed to process surgical videos. Its development was grounded in principles of transparency and long-term adaptability. This strategy ensures that the system can evolve alongside advancements in video processing, AI-based detection of sensitive visual information, and regulatory requirements. Design decisions emphasized flexibility in both deployment and use. The software architecture was developed to accommodate variable hardware capabilities, institutional constraints, and user expertise levels, from researchers with technical backgrounds to clinicians with minimal computing experience. Dependencies were carefully selected to maximize platform compatibility and minimize installation friction, enabling straightforward setup in typical clinical or academic environments. The project also adopted community-oriented development practices, such as Git-based version control, markdown issue tracking, and standardized code formatting, to maintain clarity and facilitate collaboration. The software is also released under a source-available non-commercial license (PolyForm Noncommercial) to ensure that the codebase remains transparent and reproducible, and can be openly reviewed and extended within the academic community.
2.3 Testing phase
A second survey was conducted to evaluate the final version of Endoshare in external clinical environments, constituting the user-oriented testing phase of the SDLC. Distributed via Google Forms, the survey targeted practicing surgeons who were given early access to Endoshare, which was installed on their personal or work computers. On top of the heuristic metrics used in the analysis phase, the design incorporated elements from the Technology Acceptance Model (TAM) [29] to assess Perceived Usefulness (PU) and Perceived Ease Of Use (PEOU) (Table 1). The survey included both closed and open-ended questions, grouped into the following domains: user demographics and clinical background, prior experience with video processing, system usability (heuristics), evaluation of specific features (merging and de-identification), adherence to usability principles (TAM), assessment of needs for surgical video management, and overall user satisfaction.
@@ 2.3 Testing phase
2.4 Outcomes and statistical analysis
Survey data were analyzed using a mixed-methods approach. Qualitative feedback was subjected to inductive thematic analysis by two independent coders. Quantitative responses were reported as mean ± standard deviation (SD). For the technical assessment, processing times were measured across three computing platforms (Machine: Mac laptop, Windows desktop, and Linux workstation), two processing configurations (Mode: fast and advanced), and three video durations (Video: 1, 30, and 60 minutes). Each condition was repeated three times, for a total of 54 benchmark runs. Processing times were summarized for each Machine × Mode × Video condition as geometric mean, which is preferred in this context because it better characterizes variability across multiple runs of the same video on the same machine. Differences between fast and advanced modes were expressed as geometric mean ratios (fast / advanced) with bootstrap 95 % confidence intervals. To examine scaling behavior, a linear regression through the origin was fitted to model processing time as a function of video duration for each Machine × Mode, yielding slope estimates (seconds per minute) and bootstrap 95 % CIs. The slope represents the processing rate, that is, how many seconds are required to process one minute of video. All analyses were conducted in Python using pandas, SciPy, and Statsmodels.
@@ 3 Results
Endoshare is a publicly available application, freely downloadable from the following website. The source code is available at this GitHub repository. Thorough documentation is provided to support peer review, reproducibility, and future improvements.
3.1 Analysis phase
The first survey was conducted by four surgeons and four computer scientists. Both groups consistently rated the heuristics highly ($4.68/5\pm 0.40$ vs. $4.03/5\pm 0.51$ for surgeons and computer scientists, respectively). Error Prevention was the item that received the highest overall score ($4.63/5\pm 0.52$), while Recognition Rather Than Recall was the lowest ($4.00/5\pm 0.93$).
Qualitative analysis of heuristic-based feedback matched the numeric evaluation of these metrics: it revealed positive remarks regarding system consistency, error prevention, and visual minimalism, with users describing the interface as ”easy to follow” and ”not distracting from work”. However, comments linked to the heuristics of visibility and recognition highlighted recurring issues with font size, unclear labels, and the lack of visual guidance, suggesting that key elements of the workflow could be made more discoverable and self-explanatory. These items were improved during the development of Endoshare.
3.2 Endoshare application
@@ 3.3 Testing phase
All participants recorded surgical videos primarily for research purposes, with secondary uses including conferences (80%), quality assurance (70%), training (60%), documentation (40%), and education (30%). The most common capture method was via USB drives connected to the endoscopic tower (80%), while alternative methods such as edge solutions or cloud upload were used by fewer respondents ($\leq$40%), as reported in Figure 3A. Long-term storage was most frequently on external hard drives (80%), followed by personal computers (30%) and hospital servers or cloud platforms (20%) (Figure 3B).
Participants reported intention to use Endoshare mainly for clinical and translational research (100%), followed by multicenter study preparation (60%), hospital databases (40%), conferences (30%), and education (10%). Those who will mostly use the platform include researchers (90%), residents or trainees (70%), and surgeons (70%), with less involvement from educators (20%) or nursing/OR staff (20%) (Figure 3C–D).
@@
With regard to the performance benchmarking, a total of 18 conditions (3 machines $\times$ 2 modes $\times$ 3 video lengths) were tested, each repeated three times. Geometric mean processing times increased proportionally with video duration on all platforms (Figure 5).
Fig. 5: Benchmarking Endoshare performance. Processing speed of Endoshare across three hardware configurations (macOS laptop, Windows average desktop, Linux high-end desktop) for 1-, 30-, and 60-minute videos in both fast and advanced modes. Bars indicate geometric mean processing time relative to video duration.
Across all machines and video lengths, fast mode required only 21.3% of the processing time of advanced mode (pooled geometric mean ratio =0.213=0.213, 95% CI 0.1940.1940.2350.235), corresponding to an average 78.7% reduction in total.
The linear fits confirmed a near-perfect proportional relationship between processing time and video length. Mean slope estimates were $9.22~\mathrm{s/min}$ for fast mode and $69.22~\mathrm{s/min}$ for advanced mode. Processing times varied across hardware, most notably in advanced mode for longer videos, while fast mode substantially narrowed these gaps, and the relative speed advantage of fast over advanced mode remained consistent across systems. Complete results are reported in Supplementary eTable 2.
@@ 4 Discussion
Endoshare, a publicly available and platform-independent application, systematically addresses main limitations in surgical video sharing and management. Developed through an iterative, user-centered design approach, Endoshare is an easy to use, surgeon-friendly solution to streamline surgical video use, as illustrated in Figure 6.
@@
Fig. 6: Schematic overview of the Endoshare implementation within surgical practice. Endoscopic procedures are recorded and transferred via external storage to any computer with Endoshare installed; the platform then automates merging and de-identification, facilitating use and enabling privacy-preserving sharing of videos for education, multicenter research, and laboratory analysis.
@@
The pipeline starts with the recording of videos guiding minimally invasive procedures; users can then transfer via encrypted external hard drives the videos to any computer equipped with Endoshare. With a single setup step, all videos are automatically prepared for processing and export, making them immediately available for educational or research use. By default, the application launches in fast mode to maximize throughput, while an advanced interface offers power users fine‐grained control over encoding parameters, such as resolution and frame rate, when needed. A step‐by‐step user guide is bundled with the software to facilitate initial installation and setup.
@@
This application was shaped and improved through two distinct evaluation phases. In the analysis phase, the initial usability assessment guided early refinements of Endoshare, particularly in visibility, recognition, and labeling. In the subsequent testing phase, external users confirmed the strengths of the application, especially in error prevention, interface consistency, and minimalistic design, which were reflected in the high heuristic usability scores. In addition, high PU and PEOU scores, together with strong recommendation likelihood, confirmed high user acceptance. Complementing these usability findings, the technical evaluation confirmed Endoshare’s consistent performance across different hardware configurations and video durations. Processing time scaled proportionally with video length, and the system maintained stable operation on all tested machines, demonstrating robustness and cross-platform compatibility.
@@
Endoshare is characterized by an open architecture, minimal hardware dependencies, and ease of deployment. To the best of our knowledge, it represents the first freely available application in this domain, whereas comparable systems are generally integrated within proprietary commercial packages or distributed at cost. Endoshare can be seamlessly installed on any Windows, macOS, or Linux workstation and operates using standard file-system protocols. Its modular design permits rapid inclusion of new codecs or de-identification algorithms without altering core infrastructure, and its reliance on widely used libraries (PyQt5, FFmpeg, TensorFlow) ensures transparency and reproducibility. In contrast, established clinical video systems often impose licensing fees, rigid workflows, and limited customization, which can hinder adoption, particularly in resource-constrained or research-focused settings. By providing a fully auditable codebase and comprehensive user documentation, Endoshare lowers the barrier for centers to implement video sharing while keeping data management fully within the institution.
@@ 4.1 Future Developments
Functionality will also be extended to include manual review and correction of frames misclassified by the automated de-identification algorithm. A dedicated interface will allow users to inspect flagged scenes, adjust classifications, and reinitiate the de-identification process as needed. This manual override mechanism is critical for addressing edge cases and improving the overall reliability and trustworthiness of the platform in clinical settings.